Scientific Method, Statistical Method, and the Speed of Light.

R.J. Mackay and R.W. Oldford

Working Paper 2000-02 Department of Statistics and Actuarial Science University of Waterloo

ABSTRACT

What is “statistical method”? Is it the same as “scientific method”? This paper answers the first question by specifying the elements and procedures common to all statistical investigations and organizing these into a single structure. This structure is illustrated by careful examination of the first scientific study on the speed of light carried out by A.A. Michelson in 1879. Our answer to the second question is negative. To understand this a history on the speed of light up to the time of Michelson’s study is presented. The larger history and the details of a single study allow us to place the method of statistics within the larger context of science.

Keywords: statistical method, scientific method, speed of light, philosophy of science, history of science.